Consider the parametric curve described by r(t) =〈2 sin(3t), √4t, 2cos(3t)〉. Find the unit tangent vector of r(t).
1
Expert's answer
2020-05-27T18:47:11-0400

We have given the parametric curve (vector valued) function as

"r(t)=\u27e82sin(3t), \n\\sqrt{4t}\n\u200b\t\n ,2cos(3t)\u27e9".

Since, we know that tangent vector to "r(t)" is given by "\frac{dr(t)}{dt}=r'(t)" .

Thus,

"r'(t)=\big \lt \frac{d(2\sin(3t))}{dt},\frac{d(\sqrt{4t})}{dt},\frac{d(2\cos(3t))}{dt}\big \gt \\ \implies r'(t)=\big \lt 6\cos(3t),\frac{1}{\sqrt{t}},-6\sin(3t)\big \gt" .

We also know that, unit vector for any vector "v" is "\hat{v}=\frac{v}{||v||}" where, ∣∣"v"∣∣ is the norm of "v". Hence norm of tangent vector of "r(t)"  will be

"||r'(t)||=\sqrt{(6\sin(3t))^2+(\frac{1}{\sqrt{t}})^2+(-6\cos(3t))^2} \\ \implies ||r'(t)||=\sqrt{36+\frac{1}{t}}"

Therefore,

"\widehat{r'(t)}=\frac{r'(t)}{||r'(t)||}\\ \implies \widehat{r'(t)}=\big \lt \frac{6\cos(3t)}{\sqrt{36+\frac{1}{t}}},\frac{1}{\sqrt{t}\sqrt{36+\frac{1}{t}}},\frac{-6\sin(3t)}{\sqrt{36+\frac{1}{t}}}\big \gt \\ \implies \widehat{r'(t)}=\big \lt \frac{6\cos(3t)}{\sqrt{36+\frac{1}{t}}},\frac{1}{\sqrt{36t+1}},\frac{-6\sin(3t)}{\sqrt{36+\frac{1}{t}}}\big \gt"


Hence the required unit vector is "\big \lt \frac{6\cos(3t)}{\sqrt{36+\frac{1}{t}}},\frac{1}{\sqrt{36t+1}},\frac{-6\sin(3t)}{\sqrt{36+\frac{1}{t}}}\big \gt".